1. a) A piston–cylinder assembly contains air, initially at 1.9 bar, 295 K, and a volume...

1.

a)

A piston–cylinder assembly contains air, initially at 1.9 bar, 295 K, and a volume of 0.6 m3. The air undergoes a process to a state where the pressure is 1 bar, during which the pressure–volume relationship is pV = constant. Assuming ideal gas behavior for the air, determine the mass in kg.

b)

Argon contained in a closed, rigid tank, initially at 51.1°C, 2.1 bar, and a volume of 2.9 m3, is heated to a final pressure of 7.1 bar. Assuming the ideal gas model with k = 1.51 for the argon, determine the heat transfer, in kJ.

Can you plis solved and show me step by step this problem. Im having a hard time with this topic, thank you! If little notes can be provided or explanation can be given I would be really grateful, I really want to learn how to do it.
Answer #1

1. Page ① all formulas at the o end of I will provide the solution. - 0@ system Initial conditions libar = 100 kPa) Pi= 1.9 bPage ① Augon nos 1.51 Given Initial conditions Pi= 2.1 bara 210 kPa V = 2.9 m3 T1 = 5ijºc=51.1+273 T = 324.1 ķ final conditioPage ① i Heat transter 110 XL Q=1.4-151] [1210X27.9)-1 (1.4-1 J L 101-1 = 60.275) *(-607.333) Q 2167.01 les Heat ofequüved toEQ < : Scanned by CamScanner NOTE! - wolie done by the system = tve woolle done on the system a ve # cloned system codlk ( |EQ < : Ler * clockwise - wolk prodering device Anti-clock wire =) wolk absorbing device Scanned by CamScanner procen! Heat -